# Dislocation-mediated short-range order evolution during thermomechanical processing

### v1

@@ 3.5. Process-space extrapolation of steady-state SRO
Figure 5: Steady-state SRO during manufacturing processes. a) The variation of steady-state SRO $D^{\infty}_{\text{sro}}$ induced by thermomechanical processing as a function of temperature and strain rate. The dashed lines show the empirical model (eq. 7) fit to the data. b) Equilibrium SRO states obtained from annealing at different temperatures. The equivalent temperature, $T_{\text{eq}}$, is determined by matching $D^{\infty}_{\text{sro}}$ obtained from thermomechanical processing with its corresponding equilibrium SRO value. An example match is highlighted by the magenta circled markers; the corresponding $T_{\text{eq}}$ is indicated by the vertical dashed line in the inset. c) Map of $T_{\text{eq}}$ across a wide range of temperatures and strain rates, spanning three relevant manufacturing process categories. The dashed line segments indicate extrapolated values obtained from the empirical model.
The processing conditions investigated here correspond to high strain-rate deformation, characteristic of processes such as laser-induced shock or ablation[68, 69]. Nevertheless, our results can be extrapolated to predict steady-state SRO ($D^{\infty}_{\text{sro}}$) in lower strain-rate regimes more relevant to conventional materials processing. To enable such extrapolation we introduce the following empirical model:
@@
Each parameter in this model has a well-defined physical interpretation that will be discussed next. As shown in fig. 5a, at the low-temperature limit, $D^{\infty}_{\text{sro}}$ reaches a finite minimum, indicating the persistence of steady-state SRO even in the absence of any thermal activation. We refer to this residual SRO as athermal SRO $D^{\text{ath}}_{\text{sro}}$. At the high-temperature end of fig. 5a, $D^{\infty}_{\text{sro}}$ approaches a saturation value that we call saturation SRO $D^{\text{sat}}_{\text{sro}}$. The transition temperature marking the onset of rapid SRO increase with temperature is denoted as $T_{\text{o}}$, while the parameter $w$ governs the sharpness of this transition. All model parameters, except for $w$, exhibit a linear dependence on the logarithm of strain rate. Despite its simplicity, this empirical model provides an excellent fit to the simulation data (fig. 5a), enabling valid extrapolation to a broader range of processing conditions.
@@
Using the empirical model (eq. 7), we extrapolate $D^{\infty}_{\text{sro}}$ to slower strain rates and construct a corresponding map of $T_{\text{eq}}$ across a wide range of processing conditions in fig. 5c. The extrapolated strain-rate regimes are illustrated by three representative manufacturing process categories: forging/rolling[70, 71], machining[72, 73], and laser-induced shock/ablation[68, 69]. Figure 5c demonstrates that each of these categories can access a broad spectrum of SRO states, including states with $T_{\text{eq}}$ well below the melting point. For example, under a strain rate typical of forging/rolling ($20\,\text{s}^{-1}$) at a deformation temperature of $1100\,\text{K}$, the TiTaVW alloy can reach a steady-state SRO comparable in magnitude to the SRO state obtained by annealing at $2{,}200\,\text{K}$. However, as discussed in section 3.2, the chemical nature of these two SRO states would be fundamentally different.
It is important to note that our simulations were conducted at high strain rates, where diffusion-based mechanisms such as dislocation climb and point-defect migration are largely suppressed. Among these, vacancy migration is known to be responsible for SRO formation during equilibrium processes. As a result, the empirical model in eq. 7 will not capture the full contribution such diffusion-mediated mechanisms. At lower strain rates, longer timescales will allow these mechanisms to become active[74], potentially enhancing SRO further and consequently lowering $T_{\text{eq}}$ beyond the predictions of our model. Thus, the $T_{\text{eq}}$ values shown in fig. 5c, particularly at low strain rates, should be interpreted as conservative upper bounds estimates. In practice, even higher degrees of SRO may be attainable through TMP. This underscores the potential of TMP as a practical route for achieving targeted SRO states.
@@ 4. Conclusion
Building on our previous finding that dislocation motion both creates and destroys SRO, driving alloys into far-from-equilibrium states[1], we demonstrate here that during TMP the evolution of SRO follows a quantitative balance between creation and annihilation rates, $\Gamma$ and $\lambda$, and we quantify how this balance varies with temperature and strain rate. We further explain this dependence through physical features of the dislocation network — dislocation density, edge-to-screw ratio, and the population/annealing of cross-kink-induced defects — which together govern the degree of chemically biased versus mechanically biased dislocation motion.
@@
The steady-state SRO states reached under TMP are far-from-equilibrium[1], lying outside the range accessible by thermal annealing and thus expanding the accessible SRO landscape. The framework established here — linking processing parameters to $\Gamma$ and $\lambda$ through measurable dislocation properties — provides a physically grounded, transferable basis for understanding and predicting SRO evolution in chemically complex alloys. While the present results are derived from high–strain-rate atomistic simulations of a refractory high-entropy alloy, the mechanistic connections identified between dislocation density, character, and chemical bias are general, offering a route to integrate SRO evolution into broader models of microstructural development under TMP. Perhaps more importantly, the mechanistic basis for SRO evolution derived here is testable with diffraction, dislocation analysis, and chemical-correlation measurements during controlled TMP.
@@
The software for SRO quantification can be found in our ChemicalMotifIdentifier Python package[75]. The machine learning potential can be found in our Machine-learning-potentials-for-modeling-alloys- across-compositions GitHub repository[76]. Our figure style is implemented in LovelyPlots[77] under the paper style. Any custom code or data that is not currently available in these repositories can be subsequently added upon reasonable request to the corresponding author.

### v2

@@ 3.5. Process-space extrapolation of steady-state SRO
Figure 5: Steady-state SRO during manufacturing processes. a) The variation of steady-state SRO $D^{\infty}_{\text{sro}}$ induced by thermomechanical processing as a function of temperature and strain rate. The dashed lines show the empirical model (eq. 7) fit to the $T$-$D^{\infty}_{\text{sro}}$ data for each strain rate. b) Equilibrium SRO states obtained from annealing at different temperatures. The equivalent temperature, $T_{\text{eq}}$, is determined by matching $D^{\infty}_{\text{sro}}$ obtained from thermomechanical processing with its corresponding equilibrium SRO value. An example match is highlighted by the magenta circled markers; the corresponding $T_{\text{eq}}$ is indicated by the vertical dashed line in the inset. c) Map of $T_{\text{eq}}$ across a wide range of temperatures and strain rates, spanning three relevant manufacturing process categories. The dashed line segments indicate extrapolated values obtained from the empirical model fitted using $D^{\infty}_{\text{sro}}$ across all temperature and strain rate cases (see Supplementary Information section 5)
The processing conditions investigated here correspond to high strain-rate deformation, characteristic of processes such as laser-induced shock or ablation[righi_towards_2021, tang_quantifying_2024]. Nevertheless, our results can be extrapolated to predict steady-state SRO ($D^{\infty}_{\text{sro}}$) in lower strain-rate regimes more relevant to conventional materials processing. To enable such extrapolation we introduce the following empirical model:
@@
Each parameter in this model has a well-defined physical interpretation that will be discussed next. As shown in fig. 5a, at the low-temperature limit, $D^{\infty}_{\text{sro}}$ reaches a finite minimum, indicating the persistence of steady-state SRO even in the absence of any thermal activation. We refer to this residual SRO as athermal SRO $D^{\text{ath}}_{\text{sro}}$. At the high-temperature end of fig. 5a, $D^{\infty}_{\text{sro}}$ approaches a saturation value that we call saturation SRO $D^{\text{sat}}_{\text{sro}}$. The transition temperature marking the onset of rapid SRO increase with temperature is denoted as $T_{\text{o}}$, while the parameter $w$ governs the sharpness of this transition. All model parameters, except for $w$, exhibit a linear dependence on the logarithm of strain rate (details in Supplementary Information section 5). Despite its simplicity, this empirical model provides an excellent fit to the simulation data (fig. 5a), enabling valid extrapolation to a broader range of processing conditions.
@@
Using the empirical model (eq. 7), we extrapolate $D^{\infty}_{\text{sro}}$ to slower strain rates and construct a corresponding map of $T_{\text{eq}}$ across a wide range of processing conditions in fig. 5c. The extrapolated strain-rate regimes are illustrated by three representative manufacturing process categories: forging/rolling[switzner_effect_2010, nwachukwu_effects_2017], machining[jaspers_material_2002, zhang_strain_2020], and laser-induced shock/ablation[righi_towards_2021, tang_quantifying_2024]. Figure 5c demonstrates that each of these categories can access a broad spectrum of SRO states, including states with $T_{\text{eq}}$ well below the melting point. For example, under a strain rate typical of forging/rolling ($20\,\text{s}^{-1}$) at a deformation temperature of $1100\,\text{K}$, the TiTaVW alloy can reach a steady-state SRO comparable in magnitude to the SRO state obtained by annealing at $2{,}200\,\text{K}$. However, as discussed in section 3.2, the chemical nature of these two SRO states would be fundamentally different.
It is important to note that our simulations were conducted at high strain rates, where diffusion-based mechanisms such as dislocation climb and point-defect migration are largely suppressed. Among these, vacancy migration is known to be responsible for SRO formation during equilibrium processes. As a result, the empirical model in eq. 7 will not capture the full contribution of such diffusion-mediated mechanisms. At lower strain rates, longer timescales will allow these mechanisms to become active[xu_influence_2023], potentially enhancing SRO further and consequently lowering $T_{\text{eq}}$ beyond the predictions of our model. Thus, the $T_{\text{eq}}$ values shown in fig. 5c, particularly at low strain rates, should be interpreted as conservative upper bounds estimates. In practice, even higher degrees of SRO may be attainable through TMP, which underscores the potential of TMP as a practical route for achieving targeted SRO states.
The presence of diffusion-based mechanisms at lower strain rates is expected to influence the parameters in eq. 7. Since the athermal SRO ($D^{\text{ath}}_{\text{sro}}$) represents the residual SRO at low temperature, this term is not expected to change appreciably once diffusion becomes active. In contrast, the saturation SRO ($D^{\text{sat}}_{\text{sro}}$) may increase as enhanced diffusion facilitates additional SRO formation on top of dislocation-mediated processes[xu_influence_2023]. The transition sharpness parameter, $w$, would likely become larger, reflecting the stronger Arrhenius-type sensitivity of diffusion-mediated SRO formation to temperature, whereas the onset temperature $T_{\text{o}}$ should remain largely unchanged.
Finally, the empirical model parameters introduced in eq. 7, as well as their dependence on $\dot{\epsilon}$ (see Supplementary Information fig. 6), are inherently material dependent. Both the thermodynamic driving force and the kinetics of SRO are governed by atomic size mismatch, electronegativity differences, and mixing enthalpy, all of which are expected to affect the model parameters. However, SRO evolution in the presence of both dislocation motion and diffusion is a complex phenomenon, and without a systematic exploration of these effects across different alloy systems and phases, it is difficult to anticipate their collective effect. Thus, a more comprehensive investigation of these dependencies would be a valuable direction for future work.
@@ 4. Conclusion
Building on our previous finding that dislocation motion both creates and destroys SRO, driving alloys into far-from-equilibrium states[islam_nonequilibrium_2025], we demonstrate here that during TMP the evolution of SRO follows a quantitative balance between creation and annihilation rates, $\Gamma$ and $\lambda$, and we quantify how this balance varies with temperature and strain rate. We further explain this dependence through physical features of the dislocation network — dislocation density, edge-to-screw ratio, and the population/annealing of cross-kink-induced defects — which together govern the degree of chemically biased versus mechanically biased dislocation motion.
@@
The steady-state SRO states reached under TMP are far-from-equilibrium[islam_nonequilibrium_2025], lying outside the range accessible by thermal annealing and thus expanding the accessible SRO landscape. The framework established here — linking processing parameters to $\Gamma$ and $\lambda$ through measurable dislocation properties — provides a physically grounded, transferable basis for understanding and predicting SRO evolution in chemically complex alloys. A key limitation of the present study is that the results are derived from high strain rates atomistic simulations, which affects the direct applicability of the findings to lower, more practical deformation rates. However the mechanistic connections identified between dislocation density, character, and chemical bias are general, offering a route to integrate SRO evolution into broader models of microstructural development under TMP. Perhaps more importantly, the mechanistic basis for SRO evolution derived here is testable with diffraction, dislocation analysis, and chemical-correlation measurements during controlled TMP.
@@
The software for SRO quantification can be found in our ChemicalMotifIdentifier Python package[cmi_github]. The machine learning potential can be found in our Machine-learning-potentials-for-modeling-alloys- across-compositions GitHub repository[gitrepo]. Our figure style is implemented in LovelyPlots[lovelyplots] under the paper style. Any custom code or data that is not currently available in these repositories can be subsequently added upon reasonable request to the corresponding author.